[Ncep.nhc.nco_contacts] TSB Morning Rounds - Tue Jul 05, 2016

TSB Admin nhc.tsbadmin at noaa.gov
Tue Jul 5 14:18:46 UTC 2016


Here is today's summary of NHC computer operations:
--- Craig Mattocks

----------

TAFB:
1. N-AWIPS/AWIPS-2: METEOSAT pseudo-color satellite products have not been  
ingesting since about 0430Z 02 July. TSB N-AWIPS expert Matt Sardi and TSB  
AWIPS-2 expert Chris Mello will investigate.

2. GFE/NCP: Altimeter data very limited. Only SGWH2 is displayed on  
N-AWIPS. TSB N-AWIPS expert Matt Sardi will investigate.

3. The "PK" ink cartridge on the large TAFB plotter is empty and needs to  
be replaced. An NCO Help Desk ticket was submitted.
HSU:
1. ATCF: The Pacific desk ATCF workstation froze up temporarily. Rebooting  
it from the login menu failed, so the machine was power cycled. After  
reboot, the workstation remained sluggish, but usable. An urgent (1  
business day) NCO Help Desk ticket was submitted.

2. ATCF: The system time on both the Atlantic and Pacific desk ATCF  
workstations has reverted to EDT instead of UTC. This is often caused by an  
update or patch to glibc (the GNU C library). The file (or symbolic link)  
/etc/localtime needs to be reset to /usr/share/zoneinfo/UTC. An urgent NCO  
Help Desk ticket was submitted.

3. ATCF: The 00Z Dvorak fixes for EP94 from TAFB and SAB did not import  
until well after 00Z the last couple of nights. ATCF dataflow expert Monica  
Bozeman will investigate.

4. The ATCF did not populate microwave imagery for July 4th. The soure of  
this imagery is the NRL TC webpage:

http://www.nrlmry.navy.mil/TC.html

which didn't have any new microwave imagery available at that time.
Other items:
1. Yesterday's 18Z gfs_dump_alert job failed due to low AMDAR (commercial  
aircraft) data counts. AMDAR data counts were 40 percent of the 30-day  
average (774 out of 1934) for the 18Z run.

2. Data from some recent (eg, 0936 UTC, July 3) NOAA-19 polar oribiting  
satellite passes have been lost due to ground system problems at the  
Wallops Command and Data Aquisition (CDA) station.
NHC RFCs:
RFC #2082 - ETSS (Extra Tropical Storm Surge) Data Transfer to NHC - Modify  
the DBNet tables on WCOSS to send ETSS Model data to the National Hurricane  
Center. This is at the request of NHC to support operations. To be  
implemented on Thursday, July 7 at 1600 UTC.

-------------- next part --------------
An HTML attachment was scrubbed...
URL: https://www.lstsrv.ncep.noaa.gov/pipermail/ncep.nhc.nco_contacts/attachments/20160705/dfe5a848/attachment.html 


More information about the Ncep.nhc.nco_contacts mailing list